# Pairwise Independence of Representation, Classification, and Composition in Finite Extensional Magmas

### 27 Mar 2026 v1

@@ Abstract.
We identify three properties of finite extensional 2-pointed magmas: self-representation (R), the classifier dichotomy (D), and the Internal Composition Property (H). We prove they are pairwise independent. Six Lean-verified finite counterexamples at sizes $5$ through $10$ establish all six non-implications. The minimum coexistence witness has N=5N\!=\!5, which is optimal: ICP requires 3 pairwise distinct core elements, so $N\geq 5$. The three-category decomposition induced by D is an isomorphism invariant, and the ICP is logically equivalent to the standard Compose+Inert axioms. All results are formalized in Lean 4 with zero sorry.
@@ 1. Introduction
Nontrivial combinatory algebras with $\mathbf{s}$ and $\mathbf{k}$ are necessarily infinite (Theorem 3.7). Associativity is incompatible with combining a classifier and a retraction pair in a finite extensional magma (Theorem 3.6). These two obstructions carve out a specific algebraic landscape: finite, non-associative, total. They rule out the most immediate finite instantiations of PCA (vanoosten08, ; longley15, ), TCA (bethke88, ), and Turing-object (cockett08, ) universality (by cardinality), and they exclude associative settings such as semigroups, monoids, groups, and rings from supporting the R+D pattern considered here. What algebraic structure exists in the remaining landscape?
We work in the setting of finite extensional 2-pointed magmas (a finite set with a binary operation, exactly two left-absorbers, and extensionality; the framework follows universal algebra (burris81, )). We identify three algebraic properties: self-representation (R), the existence of a section-retraction pair on the non-absorber elements (the core); the classifier dichotomy (D), a clean partition of core elements into those whose action on core lands entirely in the absorbers and those whose action avoids them entirely; and the Internal Composition Property (H), the existence of a non-trivial factorization in the left-regular representation restricted to core.
Our main result is that R, D, and H are pairwise independent: no one implies any other. Six Lean-verified finite counterexamples of sizes N=5,6,8,N\!=\!5,6,8, and $10$ establish all six non-implications; coexistence witnesses at N=5N\!=\!5 and N=6N\!=\!6 show all three can hold simultaneously. The N=5N\!=\!5 bound is optimal: ICP requires 3 pairwise distinct core elements, forcing $|S|\geq 5$. Each counterexample is a concrete Cayley table found by SAT search and independently verified in Lean 4 (moura21, ), zero sorry.
Informally, R, D, and H correspond to the three components of self-simulation: encoding, classification, and internal composition. Reflective systems in practice combine all three: Smith’s 3-Lisp (smith84, ), Friedman and Wand’s reification (friedman84, ), Amin and Rompf’s collapsed towers (amin18, ), and Brown and Palsberg’s typed self-interpreters (brown16, ) all bundle representation, classification, and composition without separating them. In the finite algebraic setting, we show that this bundling is not structurally forced; whether the independence extends to $\lambda$-calculus or typed settings is an open question (Section 9).
@@
    An independence theorem: R, D, and H are pairwise independent, with Lean-verified counterexamples in all six directions at N=5N\!=\!5 through N=10N\!=\!10 (Section 6), and coexistence witnesses at N=5N\!=\!5 and N=6N\!=\!6 (Section 7).
@@ 2. Definitions
The dichotomy creates a three-category decomposition $S=Z\sqcup C\sqcup N$ where $Z=\{z_{1},z_{2}\}$ (absorbers), $C$ (classifiers), and $N$ (non-classifiers). No element can be partially classifying. We call elements in $\{z_{1},z_{2}\}$ absorber-valued (abbreviated boolean-valued when the absorbers play the role of truth values).
@@ 5. ICP Characterization
Enrichments (not used in this paper).
Two additional axioms connect the capabilities and are used in the accompanying artifact: Branch ($\rho\cdot x=f\cdot x$ when $\tau\cdot x=z_{1}$, else $\rho\cdot x=g\cdot x$, for distinguished $f,\rho\in\mathrm{core}(S)$), which bridges D and H by making the classifier’s judgment control evaluation dispatch; and a fixed-point axiom ($Y\cdot\rho=\rho\cdot(Y\cdot\rho)$ with $Y\cdot\rho\in\mathrm{core}(S)$), which extends H across the decidability boundary. Neither is required for the independence, invariance, or obstruction results of this paper.
@@ 6. Independence
    $D\not\Rightarrow H$: An extensional 2-pointed magma of size $10$ satisfying both R and D has no triple satisfying the ICP.
@@
    $H\not\Rightarrow D$: An extensional 2-pointed magma of size $10$ satisfies ICP but violates the classifier dichotomy.
@@
    $D\not\Rightarrow R$: An extensional 2-pointed magma of size $5$ satisfies the classifier dichotomy but admits no retraction pair.
@@
    $H\not\Rightarrow R$: An extensional 2-pointed magma of size $6$ satisfies ICP but admits no retraction pair.
@@
$R\not\Rightarrow D$ (N=8N\!=\!8, the Countermodel).
An $8\times 8$ Cayley table (Appendix A) defining an FRM with $s=2$, $r=3$. The retraction pair works: $r\cdot(s\cdot x)=x$ for all core $x\in\{2,\ldots,7\}$ (e.g., $3\cdot(2\cdot 5)=3\cdot 6=5$). Element 4 is a classifier: its row on core is $(0,0,0,0,1,0)\subseteq\{z_{1},z_{2}\}$. But element 5’s row on core is $(6,2,1,1,1,1)$: it maps core inputs 2 and 3 to non-absorber values (6 and 2), while mapping inputs 4–7 to absorber value 1. This mixed behavior violates the classifier dichotomy: element 5 is neither entirely absorber-valued nor entirely non-absorber-valued on core. Self-representation holds; self-description fails. Lean file: Countermodel.lean.
@@
A $6\times 6$ Cayley table (Appendix A) defining an extensional 2-pointed magma with a retraction pair ($s=r=2$, an involution on core) but no ICP. The core has 4 elements, more than the 3 pairwise distinct witnesses ICP requires, yet no triple satisfies the ICP. Lean proves $\neg\,\texttt{HasICP}$ by exhaustive search via decide. Lean file: E2PM.lean, theorem s_not_implies_icp_structural.
$D\not\Rightarrow H$ (N=10N\!=\!10).
A $10\times 10$ Cayley table satisfying D (and also R). Lean proves that no triple $(a,b,c)$ of pairwise distinct non-absorbers satisfies the ICP, by exhaustive search via native_decide. Lean files: Countermodels10.lean, ICP.lean (theorem dNotH_no_icp).
$H\not\Rightarrow D$ (N=10N\!=\!10).
A $10\times 10$ Cayley table defining an extensional 2-pointed magma with a retraction pair, satisfying ICP (witnessed by a=8,b=6,c=7a\!=\!8,b\!=\!6,c\!=\!7), yet violating the dichotomy: element 5 has both absorber-valued and non-absorber-valued outputs on core. Lean file: Countermodels10.lean.
$D\not\Rightarrow R$ (N=5N\!=\!5).
A $5\times 5$ Cayley table defining an extensional 2-pointed magma with the classifier dichotomy (classifier at index 2, non-classifiers at 3 and 4), but admitting no retraction pair. Lean proves $\neg\,\texttt{HasRetractPair}$ by exhaustive search over all candidate pairs via decide. Lean file: E2PM.lean, theorem d_not_implies_s.
$H\not\Rightarrow R$ (N=6N\!=\!6).
A $6\times 6$ Cayley table defining an extensional 2-pointed magma satisfying ICP (witnessed by a=2,b=4,c=5a\!=\!2,b\!=\!4,c\!=\!5), but admitting no retraction pair. Lean file: E2PM.lean, theorem h_not_implies_s.
@@
Methodology.
Each counterexample was found by encoding the relevant capability axioms (and the negation of the target capability) as a SAT problem over an $n\times n$ Cayley table, then solving with Z3. The search space is $n^{n^{2}}$ (all possible binary operations on $\mathrm{Fin}(n)$); the axioms reduce this to a manageable constraint system. The sizes N=8N\!=\!8 and N=10N\!=\!10 were discovered by incrementing $n$ from the minimum until SAT returned a model. Once found, each table was frozen in counterexamples.json and independently verified in Lean. The counterexample sizes are not claimed to be universally minimal: they are minimal among the role assignments tested (SAT returned UNSAT at all smaller sizes with the same role assignments), but different role assignments or overlapping encodings could in principle yield smaller witnesses. We state no lower bounds on counterexample size.
@@ 7. Coexistence
There is a witness at N=10N\!=\!10 in which the principal R+D+H roles (2 absorbers, 2 retraction pair, 1 classifier, 3 ICP witnesses) occupy 8 distinct elements, with 2 additional elements serving the enrichment axioms (Branch and Y, defined in Section 5).
@@
A concrete $10\times 10$ Cayley table (Appendix A) satisfies all three capabilities with 10 distinct role assignments. Additionally, the enrichments Branch and Y hold. Lean file: Witness10.lean. ∎
@@ 8. Related Work
Reflective towers.
Smith’s 3-Lisp (smith84, ) introduced the reflective tower, an infinite chain of meta-interpreters. Subsequent work (friedman84, ; danvy88, ) explored reification, intensions, and live meta-modification in reflective towers. Amin and Rompf (amin18, ) showed how to compile through the tower via stage polymorphism. All of these systems combine self-representation, self-description, and self-execution without separating them. Our contribution is to show that in finite extensional 2-pointed magmas, the three capabilities are separable: no capability implies any other.
@@ 9. Discussion
R, D, and H are three algebraically simple and independently motivated properties in the landscape carved out by the obstructions. K-infinity and no-associativity rule out all classical settings; what remains are finite non-associative extensional magmas with absorbers. In this setting, the three basic structural questions are: can the algebra encode its own elements (R)? Can elements be cleanly partitioned by behavioral type (D)? Does the left-regular representation contain a non-trivial factorization (H)? The independence theorem confirms these are genuinely three questions, not one. Each property also connects to a standard algebraic or categorical concept. Capability R is a section-retraction pair (maclane71, ): two core elements $s,r$ satisfying $r\circ s=\mathrm{id}$ on core. Capability D is a dichotomic purity condition on the left-regular representation: every core element’s action on core is either fully absorber-valued or fully non-absorber-valued, with no partial classification. This isolates the Boolean aspect of a subobject classifier (johnstone02, ): classification takes values in a two-element set with no intermediate truth values. It does not capture the universal aspect. In concrete witnesses, classifiers classify a fixed partition of core (2 of $2^{|\mathrm{core}|}$ subsets), not arbitrary subobjects, and the left-multiplication maps are not closed under composition, so the categorical framework in which to state a universal property does not arise. D is a separation condition on the morphism space, not a universal property. Capability H asks for a non-trivial factorization in the left-regular representation $\{L_{a}\}$. The no-associativity obstruction (Theorem 3.6) predicts that associative categorical frameworks will not apply to either connection; this is confirmed by the composition-closure failure noted above. Whether D corresponds to a known concept in non-associative algebra, or constitutes a genuinely new structural property, remains open. Whether H is equivalent to a standard categorical universal property also remains open.
@@
Self-simulation.
R, D, and H can be viewed as algebraically separable ingredients in one natural architecture for self-simulation (computing an algebra’s own operation table from within): R provides encoding, D provides classification, and H provides internal composition. Formalizing non-trivial self-simulation in finite total magmas is an open problem. Non-trivial self-simulation in the PCA (vanoosten08, ) or Turing category (cockett08, ) sense requires a fixed evaluation mechanism and universality over an open-ended domain (brown16, ). Finite total magmas lack both: the evaluator is the magma operation itself, the domain is the $n^{2}$ table entries (closed and finite), and the undecidability that prevents triviality in infinite settings is absent. This is the same structural feature that produces the K-infinity and no-associativity obstructions. A companion artifact demonstrates a reflective tower built on a 16-element algebra satisfying R, D, H with additional computational axioms (Branch, Y-combinator); the full computational system is described elsewhere.
@@
Whether the three-capability separation appears in other settings ($\lambda$-calculus, typed systems, infinite algebras) is the primary open question. Theorem 3.6 constrains the target: any such translation must work without associativity or identity. A categorical characterization of H as a universal property (stronger than the analogy above) is also open. Two further directions connect to the frameworks discussed in Section 8: (1) the extraction theorem: do the emergent behavioral roles (absorbers, classifiers, non-classifiers) formally yield a Turing category when used as objects, with role-respecting elements as morphisms? (2) the realizability connection: does the classifier dichotomy, viewed as a property of an underlying PCA, predict specific logical principles in the associated realizability topos?
The cost of legibility.
Additional organizational axioms push the minimum coexistence size to N=12N\!=\!12 (SAT-verified) and produce interpretable role structure; these are explored in a companion paper.
Realization freedom.
Each capability admits multiple axiom forms; the independence result holds for all variants.

### 04 Apr 2026 v2

@@ Abstract.
We identify three properties of finite extensional 2-pointed magmas: self-representation (R), the classifier dichotomy (D), and the Internal Composition Property (H). We prove they are pairwise independent. Lean-verified finite counterexamples at sizes $4$ through $10$ establish all six non-implications, four with provably tight bounds. The minimum coexistence witness has N=5N\!=\!5, which is optimal: ICP requires 3 pairwise distinct core elements, so $N\geq 5$. The three-category decomposition induced by D is an isomorphism invariant, and the ICP is logically equivalent to the standard Compose+Inert axioms. All results are formalized in Lean 4 with zero sorry.
@@ 1. Introduction
A finite magma is the Cayley table of a binary operation, stripped of syntax and types. Reflective systems combine encoding, classification, and composition (smith84, ; brown16, ). Are these algebraically independent, or does one force another? We study this question in finite extensional 2-pointed magmas (a finite set with a binary operation, exactly two left-absorbers, and extensionality; the framework follows universal algebra (burris81, )).
Two obstructions confirm this is the right setting. Nontrivial combinatory algebras with $\mathbf{s}$ and $\mathbf{k}$ are necessarily infinite (Theorem 3.7), ruling out the most direct finite analogues of PCA/TCA-style (vanoosten08, ; longley15, ; bethke88, ) and Turing-object-style (cockett08, ) universality. Associativity is incompatible with combining a classifier and a retraction pair (Theorem 3.6), excluding semigroups, monoids, groups, and rings. Finite extensional 2-pointed magmas are what remains. We identify three algebraic properties corresponding to these components: self-representation (R), the existence of a section-retraction pair on the non-absorber elements (the core); the classifier dichotomy (D), a clean partition of core elements into those whose action on core lands entirely in the absorbers and those whose action avoids them entirely; and the Internal Composition Property (H), the existence of a non-trivial factorization in the left-regular representation restricted to core.
Our main result is that R, D, and H are pairwise independent: no one implies any other. Lean-verified finite counterexamples at sizes N=4,5,6,8,N\!=\!4,5,6,8, and $10$ establish all six non-implications; coexistence witnesses at N=5N\!=\!5 and N=6N\!=\!6 show all three can hold simultaneously. The N=5N\!=\!5 bound is optimal: ICP requires 3 pairwise distinct core elements, forcing $|S|\geq 5$. Each counterexample is a concrete Cayley table found by SAT search and independently verified in Lean 4 (moura21, ), zero sorry. Whether the independence extends to $\lambda$-calculus or typed settings is an open question (Section 9).
@@
    An independence theorem: R, D, and H are pairwise independent, with Lean-verified counterexamples in all six directions at N=4N\!=\!4 through N=10N\!=\!10 (Section 6), and coexistence witnesses at N=5N\!=\!5 and N=6N\!=\!6 (Section 7).
@@ 2. Definitions
The dichotomy creates a three-category decomposition $S=Z\sqcup C\sqcup N$ where $Z=\{z_{1},z_{2}\}$ (absorbers), $C$ (classifiers), and $N$ (non-classifiers). No element can be partially classifying. We call outputs in $\{z_{1},z_{2}\}$ absorber-valued.
@@ 4. Decomposition Invariance
Corollary 4.3 (Canonicity of the decomposition).
The $Z/C/N$ partition is the unique coarsest non-trivial isomorphism-invariant partition of the core. Any absorber-preserving isomorphism preserves each class individually (Theorem 4.2), so no coarser partition of core into two non-empty parts can be invariant unless it coincides with $C/N$. The classification is determined by a single algebraic invariant: whether an element’s row on core has image contained in $\{z_{1},z_{2}\}$.
Theorem 4.4 (Capability Invariance).
All three capabilities R, D, and H are preserved by absorber-preserving isomorphisms of extensional 2-pointed magmas. If $\varphi:M_{1}\to M_{2}$ is an absorber-preserving isomorphism and $M_{1}$ satisfies R (resp. D, H), then $M_{2}$ satisfies R (resp. D, H). The three capabilities are intrinsic algebraic properties, not presentation artifacts.
Proof.
R: if $(s,r)$ is a retraction pair in $M_{1}$, then $(\varphi(s),\varphi(r))$ is a retraction pair in $M_{2}$. The mutual-inverse equations transport through the homomorphism, and anchoring ($r\cdot z_{1}=z_{1}$) transfers by zero-preservation.
D: the classifier $\varphi(\tau)$ is absorber-valued in $M_{2}$ (by the homomorphism and zero-preservation). The global dichotomy transfers by surjectivity: every element in $M_{2}$ has a preimage whose dichotomy status transports. Non-degeneracy transfers by injectivity.
H: the ICP triple $(\varphi(a),\varphi(b),\varphi(c))$ satisfies the factorization equation in $M_{2}$. Injectivity preserves pairwise-distinctness and non-triviality; surjectivity preserves core-preservation.
Algebraic proofs (no decide); Lean file: CapabilityInvariance.lean. ∎
@@ 6. Independence
    $D\not\Rightarrow H$: An extensional 2-pointed magma of size $4$ satisfying D has only 2 core elements, so ICP is vacuously unsatisfiable (pigeonhole).
@@
    $H\not\Rightarrow D$: An extensional 2-pointed magma of size $5$ satisfies ICP but violates the classifier dichotomy. This is tight: ICP requires $N\geq 5$.
@@
    $D\not\Rightarrow R$: An extensional 2-pointed magma of size $4$ satisfies the classifier dichotomy but admits no retraction pair. This is tight: D requires $N\geq 4$.
@@
    $H\not\Rightarrow R$: An extensional 2-pointed magma of size $5$ satisfies ICP but admits no retraction pair. This is tight: ICP requires $N\geq 5$.
@@
$R\not\Rightarrow D$ (N=8N\!=\!8).
An $8\times 8$ Cayley table (Appendix A) defining an FRM with $s=2$, $r=3$. Element 4 is a classifier, but element 5’s row on core is $(6,2,1,1,1,1)$, mixing absorber and non-absorber outputs, violating the dichotomy. SAT confirms the non-implication is achievable as early as N=3N\!=\!3, but vacuously: with only 1 core element, the dichotomy cannot be formulated (it requires both a classifier and a non-classifier). The N=8N\!=\!8 witness is the smallest with enough core structure for D to be non-vacuously falsifiable. Lean file: Countermodel.lean.
@@
A $6\times 6$ Cayley table (Appendix A) defining an extensional 2-pointed magma with a retraction pair ($s=r=2$, an involution on core) but no ICP. The core has 4 elements, more than the 3 pairwise distinct witnesses ICP requires, yet no triple satisfies the ICP. As with $R\not\Rightarrow D$, the non-implication is vacuously achievable at N=3N\!=\!3 (only 1 core element), but the N=6N\!=\!6 witness is the smallest where ICP failure is structural rather than dimensional. Lean file: E2PM.lean, theorem s_not_implies_icp_structural.
$D\not\Rightarrow H$ (N=4N\!=\!4, tight).
The minimal R+D model at N=4N\!=\!4 (kripke4 in Appendix A) has only 2 core elements. ICP requires 3 pairwise distinct non-absorbers, so it is vacuously unsatisfiable. Lean file: ICP.lean, theorem kripke4_no_icp. The N=10N\!=\!10 D+R model in Countermodels10.lean complements this: it has 8 core elements (more than enough for ICP to be non-vacuous), yet no triple satisfies ICP. The formal non-implication holds at N=4N\!=\!4; the N=10N\!=\!10 witness shows it persists even when H has room to exist.
$H\not\Rightarrow D$ (N=5N\!=\!5, tight).
A $5\times 5$ Cayley table satisfying ICP but with no classifier: all three core elements map core to core (no element produces only absorber-valued outputs on core). The bound is tight: ICP requires $N\geq 5$. Lean file: E2PM.lean, theorem h_not_implies_d_tight.
$D\not\Rightarrow R$ (N=4N\!=\!4, tight).
A $4\times 4$ Cayley table with the classifier dichotomy (classifier at index 2, non-classifier at index 3) but no retraction pair among the 4 possible assignments from core. The bound is tight: D requires $N\geq 4$. Lean file: E2PM.lean, theorem d_not_implies_s_tight.
$H\not\Rightarrow R$ (N=5N\!=\!5, tight).
A $5\times 5$ Cayley table satisfying ICP but admitting no retraction pair. The bound is tight: ICP requires $N\geq 5$. Lean file: E2PM.lean, theorem h_not_implies_s_tight.
@@
Methodology and minimality.
[…]
Scaling.
The independence is not a small-size artifact. SAT search confirms that all six non-implications are satisfiable at every size from their respective minima through N=15N\!=\!15 (the solver limit). No size was found where any direction becomes UNSAT. Whether the independence holds at all sizes (for all $n\geq 5$, there exist witnesses at size $n$ for each direction) is a natural conjecture supported by this data.
@@ 7. Coexistence
There is a witness at N=10N\!=\!10 in which the principal R+D+H roles (2 absorbers, 2 retraction pair, 1 classifier, 3 ICP witnesses) occupy 8 distinct elements.
@@
A concrete $10\times 10$ Cayley table (Appendix A) satisfies all three capabilities with distinct role assignments. Lean file: Witness10.lean. A witness with $s\neq r$ (non-degenerate retraction) is exhibited at N=6N\!=\!6 (Appendix A). ∎
@@ 9. Discussion
Extensionality identifies each element with its row in the Cayley table. The elements of a finite extensional 2-pointed magma are therefore a family of typed maps $\{L_{a}^{\mathrm{core}}:\mathrm{core}\to\mathrm{core}\cup Z\}$, one per element, with the absorber maps $L_{z_{1}},L_{z_{2}}$ being constant. R, D, and H correspond to the three fundamental structural questions about such a family:
  • Splitting (R): does the identity on core split through the family? That is, do some $L_{s},L_{r}$ satisfy $L_{r}\circ L_{s}=\mathrm{id}_{\mathrm{core}}$? This is a section-retraction pair (maclane71, ).
  • Codomain purity (D): does each map hit $\mathrm{core}$ or $Z$, never both? This is the unique coarsest non-trivial isomorphism-invariant partition of core (Corollary 4.3).
  • Factorization (H): does the family contain a non-trivial composition, $L_{a}=L_{c}\circ L_{b}$ with $L_{b}$ core-preserving and $L_{a}$ non-constant? This is the weakest factorization condition that is not vacuously satisfiable: weakening either the pairwise-distinctness or the non-triviality requirement makes every retraction-equipped magma satisfy it trivially (Lean-verified separations in ICP.lean).
Each property is minimal in its own sense. R is the standard notion of splitting. D is the coarsest invariant partition. H is the weakest non-degenerate factorization. The independence theorem confirms these are genuinely three questions, not one. The precise categorical status of D and H is discussed further in the categorical connections paragraph below; whether either corresponds to a known universal property remains open.
Categorical connections and their limits.
Several constructions in this paper have natural categorical readings. The core $\mathrm{core}(S)$ is a retract of the carrier via $(s,r)$, in the standard sense of (maclane71, ). The three-category decomposition defines a functor from the category $\mathbf{E2PM}$ (extensional 2-pointed magmas with absorber-preserving homomorphisms) to the category of three-block set partitions: Theorems 4.2 and 4.4 establish that both the decomposition and the three capabilities are preserved by morphisms. The classifier dichotomy is a codomain-purity condition on the family of left-multiplication maps $\{L_{a}^{\mathrm{core}}:\mathrm{core}\to\mathrm{core}\cup Z\}$, partitioning them by image type.
The categorical analogy breaks down at composition. The left-multiplication maps $L_{a}^{\mathrm{core}}$ are not closed under composition: $L_{a}\circ L_{b}$ need not equal $L_{c}$ for any element $c$. For instance, in the N=5N\!=\!5 coexistence witness, $L_{2}\circ L_{3}$ on core gives $(0,2,0)$, which is not the core row of any element. This means the row family does not form a transformation monoid, and standard categorical constructions (internal hom, enrichment, Turing morphism) cannot be stated internally. The no-associativity theorem (Theorem 3.6) predicts this failure: associativity of the magma operation would be a prerequisite for the left-multiplication maps to form a monoid, and the theorem shows associativity is incompatible with the R+D structure. The composition-closure failure is not a gap in the formalization but a consequence of the same obstruction that motivates the setting.
@@
Whether the three-capability separation appears in other settings ($\lambda$-calculus, typed systems, infinite algebras) is the primary open question. Theorem 3.6 constrains the target: any such translation must work without associativity or identity. Each capability admits multiple axiom forms; the independence result holds for all variants. The independence also implies that any system combining all three capabilities must introduce additional axioms beyond R, D, and H to connect them; the nature of these connecting axioms is a natural direction for further work.